# Geometric QCD II: The Confining Twistor String and Meson Spectrum

### v1

@@ 17.7 Future Directions: Glueballs, Baryons and Scattering
    Glueballs: the whole construction, starting with the loop equations to the fermion determinant at the minimal surface, to the rigid twistor string and the localization of the twistor path integral for quark amplitudes at the boundary of the moduli space, naturally generalizes to the higher $1/N_{c}$ corrections. There are some important mathematical issues to be resolved for that program to materialize. This is a natural next step for this Geometric QCD theory.
@@
    Baryons: In the large-$N_{c}$ limit, baryons appear as heavy states composed of $N_{c}$ quarks. Geometrically, this corresponds to a minimal surface with a specific topology (e.g., a "Y-junction" or a surface with multiple boundaries) [5]. Extending the twistor ansatz to surfaces with higher connectivity is a natural next step.
@@
    Scattering Amplitudes: We have focused here on the two-point function (spectrum). The four-point function (meson-meson scattering) involves a minimal surface bounded by four segments. The crossing symmetry and Regge behavior of these amplitudes should emerge from the analytic continuation of the twistor data, potentially connecting this framework to the modern "Amplituhedron" methods.
18 Conclusion
In this paper, we developed the dynamical part of the Geometric QCD program. Building on [1], where the confining dressing factor $\exp\left(-\kappa S[C]\right)$ arises from the rigid Hodge-dual additive minimal surface, we formulated a concrete large-$N_{c}$ framework in which the remaining degrees of freedom can be treated without summing over fluctuating worldsheet metrics.
The first main outcome is conceptual and structural: by working in momentum loop space, we bypass the coordinate-space contact singularities and cusp pathologies that obscure the local continuum interpretation of the loop equations. In this representation, the planar loop equation becomes an algebraic–differential equation. Furthermore, by mapping the Master Field to the infinite-dimensional Cuntz-algebra “word space” (Section 10.8), we rigorously establish that a purely one-dimensional analytic Taylor–Magnus (finite-moment) solution of the exact vector momentum-loop equation cannot exist: the expansion is consistent through $\mathcal{O}(P^{\prime 6})$ but becomes algebraically inconsistent at $\mathcal{W}^{(8)}$ (C, Table 1). This violent breakdown is a theorem-level obstruction showing that a 1D closed-loop scalar Taylor–Magnus framework lacks sufficient internal degrees of freedom to absorb the macroscopic spatial stress generated by planar QCD. This breakdown prompts the introduction of additional continuous degrees of freedom; in our construction these are the 4D boundary-twistor variables that parametrize the rigid Hodge-dual minimal surface and yield the confining twistor-string representation.
The second main outcome is dynamical: we showed that the fermionic (Majorana) worldsheet degrees of freedom of the Fermi string provide the correct algebraic mechanism for planar factorization at intersections, while the rigidity of the Hodge-dual minimal surface removes the Liouville instability of the original random-surface formulation. The resulting theory may be viewed as an analytic (confining) twistor-string description in which the reduced loop-space measure is most naturally written in twistor variables, leading to an effective holographic Liouville functional determined by holomorphic extension of boundary data.
The third main outcome concerns spectroscopy and the profound realization that the quantum path integral geometrically localizes. We initially demonstrated the computability of the boundary-twistor functional integral by evaluating it on the IAS Typhon supercluster via Complex Langevin (CLE) evolution with logarithmic reweighting. To overcome the severe topological trapping and infinite-variance $1/|u|^{3}$ Cauchy tails inherent to the complexified phase space, we introduced an unsupervised phase-information clustering algorithm. Applying this to our massive stochastic histories successfully isolated two distinct, approximately linear Regge trajectories. This numerical exploration served a crucial purpose: it proved that the Confining Twistor String natively possesses internal quantum numbers (analogous to string “daughter trajectories”) emerging directly from the multiple complex saddle points (Lefschetz thimbles) of the path integral.
However, the ultimate realization of this work goes far beyond numerical stochasticity. By analyzing the analytic structure of the Twistor String, we discovered that the physical mass spectrum is deterministically governed by Catastrophe Theory. The meson poles correspond exactly to degenerate Lefschetz thimbles—bifurcation points where the complex stability matrix of the classical twistor trajectory develops a null space. Because the integration variables are coupled holomorphic twistors, this degeneracy inherently possesses a corank of 2. As dictated by the singularity theory of multiple integrals, this complex degeneracy mathematically upgrades standard semiclassical square-root branch cuts into the exact simple poles of the meson states. Summary
To summarize, what started as a rigorous continuum formulation of planar QCD via the Makeenko–Migdal loop equations culminated in an exact analytical solution for the fundamental meson spectrum. Although our numerical investigations using the Complex Langevin approach successfully provided proof-of-concept evidence for linear Regge trajectories, the true predictive power of this framework lies in its analytical semiclassical limit, which provides exact equations for the mass spectrum beyond the WKB approximation at the non-perturbative level. It is a known phenomenon that some WKB spectral equations are valid beyond the WKB approximation, because they reflect deeper hidden symmetries leading to infinite flat valleys for the steepest descent path, making the WKB estimate an exact localization of the path integral [11]. We have shown that the $S$-matrix mass poles of the theory are naturally captured by Catastrophe Theory, corresponding to non-compact flat valleys in the complexified effective action. These flat valleys, where the steepest descent integral diverges without higher-order corrections, are driven entirely by the topological monodromies of twistor poles enclosed within the unit disk.
The number of such poles serves as a rigorously conserved topological quantum number, protected by an absolute pinch-singularity barrier at the conformal boundary. The crossing of the unit circle by a twistor pole leads to a divergent action, because the conjugate poles $z_{0}$ and $1/\bar{z}_{0}$ inevitably pinch the physical integration loop. By analytically continuing the Euclidean area action to the boundary and summing over the exact purely imaginary monodromy phases, we obtain a Bohr–Sommerfeld quantization condition. For the fundamental one-pole trajectory, this yields the exact simple mass poles and the linear Regge spectrum:
$m^{2}=\pi\sigma\left(n+\frac{1}{12}\right)$ (18.1)
This analytical result not only confirms the perfectly linear Regge slope but uniquely derives the exact open-string Lüscher intercept $(+1/12)$ directly from the 2D anomaly of the Elfin determinant on rigid minimal surface. Thus, by mapping the momentum loop-space dynamics to boundary twistors, the theory does not merely approximate the confining dynamics of planar QCD; it provides its exact algebraic geometric solution. The exact analytical classification of higher topological sectors (multi-pole configurations) and their corresponding daughter trajectories is now a fully accessible problem. Consequently, Planar QCD ceases to be a theory of stochastic quantum fluctuations; it becomes a theory of pure classical complex geometry. The extraction of the mass spectrum is reduced from a fluctuating path integral to a deterministic Generalized Eigenvalue Problem for the degenerate saddle points of the effective action.
@@ Acknowledgements
I am grateful to Nima Arkani-Hamed for the invitation to present this work as a series of lectures at the IAS Particle Physics (Pizza) Seminar, which provided a stimulating environment for discussing these results. I also thank Edward Witten for valuable discussions regarding the distinction between this confining geometry and the holographic dualities of conformal field theories.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the author used Google Gemini to fix typographical errors, identify relevant mathematical terminology, clarify the notations, and improve the overall language and style. After using this tool, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.
@@
The spectral data reported in this paper are generated by numerical simulation of the Complex Langevin equation (CLE) for the discretized Confining Twistor String spectral observable, including the reweighted estimator for $1/R(E)=\Pi(E)/\Pi^{\prime}(E)$. To enable reproduction and extension of these calculations, I have made the full Mathematica ® workflow publicly available on the Wolfram Cloud, including (i) the CLE simulator and effective-action implementation, (ii) SLURM submission scripts for running independent replicas on an HPC cluster, and (iii) the analysis tools used to pool independent runs and extract the spectrum:
  • 1.
    TwistorString.wl, [52] (main CLE simulator):
  • 2.
    run_node.m (SLURM/cluster driver) [53]
  • 3.
    VerifyTwistorHistory.nb [54] (fat tail analysis of distributions for a CLE history):
  • 4.
    TwistorInformationClustering.nb [55] (spectrum extraction and plots)
  • 5.
    GeometricQCD CLE results [56] (the full set of histories of the evolution of $R(E)$ for $240$ values of energy $E\in(2.,3.)$ ran on 64 nodes of Typhon cluster, with random initial states)
The raw per-replica output consists of node-indexed binary arrays (e.g. TwistorHistory*.mx) storing, for each sampled energy value, the whole stochastic trajectory of observable resolvent $R(E)$. These files are verified by VerifyTwistorHistory.nb, testing fat tails of distribution against theoretical prediction. Then TwistorInformationClustering.nb performs robust pooling across independent replicas, analyzes the history of the phase of the resolvent, builds the statistical distribution, clusters the data by information distance between these phase distributions and produces the final spectral curves and extracted mass tables. This public release is intended to provide a practical basis for future large-scale runs (larger ensembles, finer discretization, and improved CLE diagnostics and stabilization methods).

### v2

@@ 16.5 Quark Spin Monodromy and the Unified Regge Spectrum
    Unnatural Parity ($\pi$ and $K$): $q=0\implies n=4J$, placing the sequence in the $n\equiv 0\pmod{4}$ sector and yielding the intercept $\alpha_{\pi}(0)=-1/48$.
@@
    Natural Parity ($\rho$): $q=-2\implies n=4J-2$, placing the sequence in the $n\equiv 2\pmod{4}$ sector and yielding the intercept $\alpha_{\rho}(0)=23/48$.
As demonstrated in Fig. 15, the experimental data confirms this topological identification. Because the $K$ mesons share the unnatural parity of the pion, they belong to the identical topological sector $n=4J$, with their trajectory cleanly displaced by a positive mass-squared shift $\Delta m_{s}^{2}\approx 0.17$ GeV2 reflecting the explicit $SU(3)$ flavor symmetry breaking by the heavier strange quark. The observed variation of the slopes ($\pi\sigma_{\pi}\approx 1.38$, $\pi\sigma_{\rho}\approx 1.16$, and $\pi\sigma_{K}\approx 1.58$ GeV2) remains an open dynamical question. It may arise from the interaction between a pole in $\lambda$ and a pole in $\mu$, which would locally renormalize the effective string tension. We mention this possibility here as a direction for future exploration but retain the distinct empirical fits to maintain contact with the PDG data.
We must also address a natural phenomenological question regarding our selection of experimental candidates. A critical reader comparing our trajectories with the Particle Data Group (PDG) tables will note that we have not included all listed resonances possessing the same quantum numbers. This selection was deliberate, but it is not a mere phenomenological convenience; rather, it is strictly dictated by the scope of our analytic solution. The linear Regge trajectories plotted in Fig. 15 correspond exclusively to the absolute simplest topological sector of the theory: the exact one-pole twistor solution. Our geometric framework guarantees that the remaining, unmapped hadronic resonances in the PDG are not anomalies, but are dynamically required to populate the daughter trajectories generated by multi-pole twistor configurations. Solving the complexified effective action—specifically, solving the saddle-point equations for the moduli $R_{i},z_{i}$ of the multi-pole solutions and classifying their monodromies—constitutes a deterministic, yet computationally formidable, nonlinear optimization problem in complex geometry. Systematically extracting these higher-topology solutions and mapping them onto the complete PDG database falls beyond the foundational scope of the present paper. Instead, we present this as a rigorous, data-rich research program—an open invitation to apply the sophisticated analytic machinery of string theory directly to the rich empirical data of hadronic spectroscopy.
Figure 15: Unified Chew-Frautschi plot ($J$ vs $m^{2}$) showing the topological identification of the $\pi$ ($n=4J$), $K$ ($n=4J$) and $\rho$ (n=4J2n=4J-2) trajectories. The horizontal error bars represent propagated mass uncertainties, and labels indicate the experimental masses in MeV. The $K_{3}(2320)$ is marked with a question mark as it is currently omitted from the PDG Summary Tables pending further experimental confirmation.
16.6 The microscopic origin of the Lüscher term: elves instead of string vibrations
Our exact derivation of the strictly bounded $\pi/12$ topological phase unveils a fundamentally different, purely microscopic physical origin for the universally mandated Lüscher term, representing the absolute culmination of our theory.
In standard Effective String Theory, this famous constant is interpreted strictly via the macroscopic Casimir effect. The standard calculation dictates that one must manually quantize the zero-point fluctuations of a macroscopic, highly excited “rubber band” string bending and vibrating across the transverse spacetime dimensions. The $\pi/12R$ effective potential is then forcefully extracted from the ultraviolet divergence of these vibrating spatial standing waves via Riemann zeta regularization.
Our twistor-string formulation completely upends this paradigm. We bypassed the macroscopic quantization of target-space wiggles entirely. Instead, the quantum mechanics takes place directly on the 2-dimensional worldsheet. The Lüscher term originates intrinsically from the quantum anomaly of the microscopic zero-point fluctuations of the heavy fermions—the “elves”—that are confined to our minimal surface.
When evaluating the fermion determinant (required to reproduce the planar factorization of the MM equation), integrating out these fundamental elfin degrees of freedom generates the quantum anomaly, related to the curvature. Because these quantum elves live on a highly curved minimal surface governed by the singular twistor conformal factor $\rho$, regulating their vacuum fluctuations inherently breaks conformal symmetry, generating the Weyl/Polyakov trace anomaly. The direct mathematical consequence of this trace anomaly is the rigid addition of the Liouville kinetic term $\frac{1}{3\pi}|\nabla\rho|^{2}$ to the effective action.
Crucially, the numerical coefficient of this anomaly is strictly proportional to the central charge of the worldsheet—it literally counts the active quantum degrees of freedom of the elves. In a conventional generic 3-dimensional Helicoid, both chiral sectors dynamically fluctuate, which overcounts the anomaly (yielding an unphysical $\pi/6$). However, by enforcing the chiral (Hodge-dual) configuration and taking the parity-even arithmetic mean of the SD and ASD minimal areas, we dynamically halve the internal fluctuating degrees of freedom compared to the generic minimal surface. This exact symmetrization prescription makes the day: it mathematically halves the anomaly integral over the minimal surface, flawlessly dropping the topological phase to precisely $\pi/12$.
[…]
17 Planar QCD as a Geometric theory
The dynamical generation of S-matrix mass poles through the degeneracy of the twistor-string effective action is not an isolated mathematical curiosity; it is an infinite-dimensional realization of Exact WKB analysis, Picard-Lefschetz resurgence, and the geometry of Spectral Networks [15, 10]. The complexified path integral over the boundary twistors exhibits localization phenomena that strictly mirror the topological mechanisms governing Seiberg-Witten (SW) theory and the analytic continuation of gauge theories [55]. To establish this physical correspondence, we analyze the three fundamental geometric mechanisms driving the twistor string quantization:
17.1 Topological Barriers and Infinite-Dimensional Picard-Lefschetz Theory
By analytically continuing the twistor path integral into the complexified phase space, the convergence of the quantum amplitude is governed by middle-dimensional integration cycles, or Lefschetz thimbles [5]. As formulated by Witten in the analytic continuation of Chern-Simons theory [55], the topology of these thimbles—and their intersection pairings—changes strictly when the system crosses a Stokes line, resulting in a discontinuous jump in the physical vacuum state. In our twistor string framework, non-perturbative Stokes phenomena are rigorously suppressed by the conformal boundary geometry. If a twistor pole $\xi_{1}$ attempts to cross the unit circle $|z|=1$, it inevitably collides with its conjugate mirror $\xi_{2}=1/\overline{\xi}_{1}$ directly on the physical integration contour. Exactly at the moment of crossing, these two singularities inevitably collide directly on the physical integration contour $|z|=1$. As demonstrated by the exact bulk-to-boundary reduction of the Liouville action in E, this collision creates a classic, non-integrable pinch singularity. The integration loop becomes trapped between the colliding singularities at $z_{0}$ and $1/\bar{z}_{0}$ and cannot be deformed to avoid them. This generates a severe pinch singularity, creating an infinite action barrier. In the language of Picard-Lefschetz theory, this impenetrable boundary dynamically locks the Twistor String path integral into a strictly conserved topological sector defined by the enclosed twistor pole number $n$, forbidding unphysical Stokes jumps.
17.2 Complex Langevin, Cauchy’s Shield, and Flat Valleys
The rigorous exploration of these complex integration cycles is executed via the holomorphic gradient drift of the Complex Langevin Equation (CLE). In standard semiclassical formulations, evaluating the path integral at a degenerate saddle point requires formally isolating a zero mode and calculating the fluctuation determinant of the Hessian. In the CLE framework, this static abstraction is bypassed entirely: the path integral expectation value is simply the long-time average $\int d\tau$ over a single, noisy stochastic trajectory $\mathcal{V}(\tau)$.
At the critical resonance energies $E=E_{n}$, the twistor effective action develops a non-compact flat valley. The single CLE trajectory does not rest at a classical minimum; driven by the stochastic noise, it constantly climbs the steep transverse walls of the complexified functional space and bounces back. Simultaneously, experiencing no deterministic restoring force along the zero mode, the trajectory drifts unimpeded along the flat valley. In this representation, the macroscopic pre-exponential factor (the transition residue) is simply proportional to the total Langevin time the trajectory spends accumulating windings, completely eliminating the need for formal Hessian factorization.
A fundamental mathematical question arises: how can this violently stochastic trajectory $\mathcal{V}(\tau)$ generate a pristine, exact Bohr-Sommerfeld quantization phase without acquiring infinite quantum corrections from the nonlinear effects amplifying the noise in transverse directions?
The resolution lies in the analytic continuation of the twistor fields. The CLE directly evolves the boundary data $\mathcal{V}_{bound}(\tau)=\mathcal{V}(z,\tau)\big|_{|z|=1}$. However, the twistors defining our symplectic form $\mathcal{O}$ are analytically continued from this boundary into the bulk. If we represent the twistor field as a meromorphic function, such as $\lambda(z)=\sum_{k}\frac{R_{i}}{z-z_{i}}$, its boundary values are:
$\lambda(\theta)=\sum_{k}\frac{R_{i}}{e^{i\theta}-z_{i}}$ (17.1)
The stochastic CLE evolution simply forces these boundary parameters to evolve with Langevin time $\tau$:
$\lambda(\theta,\tau)=\sum_{k}\frac{R_{i}(\tau)}{e^{i\theta}-z_{i}(\tau)}$ (17.2)
Upon analytic continuation back into the bulk, we obtain:
$\lambda(z,\tau)=\sum_{k}\frac{R_{i}(\tau)}{z-z_{i}(\tau)}$ (17.3)
The crucial observation is that the function remains strictly meromorphic at every instant of Langevin time. We obtain moving poles with evolving residues, but the fundamental meromorphic structure is absolutely preserved. Therefore, the infinite-dimensional stochastic evolution of the boundary fields rigorously and exactly reduces to the finite-dimensional, continuous $\tau$-dependence of the moduli $\{R_{i}(\tau),z_{i}(\tau)\}$.
The infinite-dimensional phase accumulation thereby rigidly reduces to a standard one-dimensional contour integral over the physical boundary:
$\Delta S=\oint_{|z|=1}\omega(z,\tau)\,dz$ (17.4)
Here, the absolute geometric power of Cauchy’s integral theorem manifests itself. At any instant $\tau$, this contour integral is completely determined by the sum of the residues at the enclosed twistor poles.
To evaluate the full physical amplitude, the path integral must be decomposed into a discrete sum over distinct topological sectors. These sectors are labeled by the winding number $k$ of the target-space minimal surface (or equivalently, the number of times the physical boundary parameter traverses the twistor poles). The full quantum amplitude is thus a superposition of path integrals, each mathematically explored by its own Complex Langevin evolution. Because the twistor effective action is strictly additive over these cycles, the total action for the $k$-th topological sector is precisely $S_{k}=S_{eff}^{(1)}+(k-1)\Delta S_{eff}$.
In the CLE formulation for the $k$-th sector, the stochastic noise is opposed by the deterministic drift and the transverse Hessian, which now scale linearly with the topological winding number: $\mathcal{H}_{k}\to k\nabla^{2}\Delta S_{eff}$. Consequently, the residual stochastic variance in this sector is inversely proportional to the Hessian, scaling exactly as $\langle\delta\mathcal{V}^{2}\rangle\propto 1/k$. If the CLE noise were to continuously jitter the moduli $\{R_{i},z_{i}\}$ across the different sectors, the topological action increment $\Delta S_{eff}$ would randomly vary from one winding to the next. This would destroy the perfect phase coherence required for the geometric series to diverge, preventing the formation of a resonance.
However, the macroscopic S-matrix pole is analytically generated exclusively by the asymptotic tail of this superposition, where $k\to\infty$. In this asymptotic limit, the transverse restoring force becomes strictly infinite, and the stochastic variance vanishes identically. The stochastic noise is completely crushed by the macroscopic weight of the multiple windings. The moduli $\{R_{i},z_{i}\}$ rigidly freeze at their exact classical saddle-point values for all large-$k$ sectors (validating the algebraic minimization of the single-pole radius $R$ in our spectral derivation). The topological phase is thus completely shielded from quantum contamination, ensuring that $\Delta S_{eff}$ remains strictly constant. The superposition over topological sectors safely diverges as a pristine geometric series $\sim(1-\exp\left(i\Delta S_{eff}\right))^{-1}$, analytically generating the exact S-matrix pole.
18 Conclusion: The Solution of Planar QCD
In this paper, we have presented the exact analytic solution of the planar MM loop equations, thereby solving large-$N_{c}$ QCD in the continuum limit. Building on Part I, where the confining dressing factor $\exp\left(-\kappa S[C]\right)$ was shown to arise from the rigid Hodge-dual additive minimal surface, we have constructed the complete dynamics without summing over fluctuating worldsheet metrics.
Our first main outcome is structural: by working natively in momentum loop space, we bypass the coordinate-space contact singularities and cusp pathologies that obscure the local continuum limit of the MM loop equations. We rigorously established that a purely one-dimensional analytic Taylor-Magnus solution of the exact vector momentum-loop equation cannot exist. The expansion is mathematically consistent through $\mathcal{O}(P^{\prime 6})$ but violently breaks down at $\mathcal{W}^{(8)}$ due to a Fredholm rank deficiency. This topological obstruction proves that a 1D closed-loop scalar framework lacks the internal degrees of freedom to absorb the macroscopic spatial stress of Planar QCD. This anomaly natively forces the introduction of 4D continuous boundary-twistor variables to parametrize the rigid minimal surface, yielding the confining twistor-string representation.
Our second main outcome is dynamical: we demonstrated that the internal Majorana fermions (the “elves”) on the worldsheet provide the precise algebraic mechanism required for planar factorization at intersections. Crucially, the bulk Liouville term in our effective action originates microscopically from the zero-point fluctuations of these hidden fermions. The conformal anomaly of these heavy elves yields exactly the Lüscher term ($\pi/12$) in the effective string potential, cleanly generating the universally correct intercepts 1/48-1/48 and $23/48$ for the physical $\pi$ and $\rho$ Regge trajectories without any ad hoc assumptions of macroscopic string vibrations.
In a profound sense, this construction explicitly realizes Edward Witten’s original “Master Field” hypothesis—the existence of a purely classical description for the large-$N_{c}$ limit of QCD. While this Master Field does not manifest as a single classical gauge field configuration in coordinate spacetime, we have discovered its exact geometric realization in loop space. The classical geometry uncovered here identifies the underlying master object as the rigid Hodge-dual minimal surface. Consequently, the phase-space path integral over quark loops, once transformed into two-dimensional boundary twistor variables, exhibits exact geometric localization governed by Catastrophe Theory. The true Master Field of Planar QCD is thus a critical classical trajectory in twistor space—a geometric structure conceptually parallel to the spectral curves of Seiberg-Witten theory.
By establishing the Master Field as this rigid twistor geometry, our approach naturally avoids the “graviton problem” typical of AdS/CFT gravity holography. Because our bulk geometry is not dynamical, there are no propagating bulk graviton modes that would lack a physical counterpart in confining QCD. All propagating degrees of freedom are strictly confined to the boundary, while bulk quantities (such as the Liouville field) are rigidly determined by the analytic continuation of boundary data. Therefore, our framework is best understood as gauge holography, where the bulk serves as a rigid encoding of the boundary dynamics rather than gravity holography or a topological string theory.
Ultimately, Planar QCD ceases to be a theory of stochastic quantum fluctuations; it becomes a theory of pure classical complex geometry. The extraction of the exact mass spectrum reduces from a fluctuating path integral to a deterministic Generalized Eigenvalue Problem. By summing over the purely imaginary monodromy phases of the twistor poles protected by the pinch-singularity boundary barrier, we obtain an exact Bohr-Sommerfeld quantization condition. For the fundamental one-pole trajectory, this yields the unified linear Regge spectrum:
$m^{2}=\frac{\pi\sigma}{4}\left(n+\frac{1}{12}\right)$ (18.1)
where $n=4J$ for the unnatural parity ($\pi,K$) trajectories and n=4J2n=4J-2 for the natural parity ($\rho$) trajectory. This analytic solution passes flawlessly through experimental data on Figure 15. The only outlier ($3\sigma$) is $K_{3}(2320)$ which is marked as "unconfirmed" in the PDG tables. The exact mass spectrum of Planar QCD is thus deterministically governed by the degenerate saddle points of the complexified twistor-string.
18.1 Future Directions: Glueballs, Baryons, and Scattering
The representation of Planar QCD as an Analytic Twistor String opens several immediate avenues for future research:
  • 1.
    Glueballs: The entire construction, from the loop equations and the fermion determinant on the minimal surface to the localization of the twistor path integral, naturally generalizes to higher $1/N_{c}$ corrections. Resolving the mathematical specifics of this expansion is the natural next step to analytically extract the glueball spectrum.
  • 2.
@@ 18.1 Future Directions: Glueballs, Baryons, and Scattering
    Baryons: In the large-$N_{c}$ limit, baryons appear as heavy states composed of $N_{c}$ quarks. Geometrically, this corresponds to a minimal surface with a specific topology (e.g., a “Y-junction” or a surface with multiple boundaries). Extending the twistor ansatz to surfaces with higher connectivity will allow for the exact geometric quantization of baryon masses.
@@
    Scattering Amplitudes: Meson-meson scattering involves a minimal surface bounded by four segments. The crossing symmetry and Regge behavior of these amplitudes should emerge naturally from the analytic continuation of the twistor data, potentially connecting this confining geometric framework directly to modern Amplituhedron methods.
@@ Acknowledgements
I wish to express my deepest gratitude to Nima Arkani-Hamed for his hospitality and for the invitation to present the complete derivation of this solution in a series of lectures at the IAS Particle Physics (Pizza) Seminar. The intellectual rigor and stimulating environment he fosters at the Institute were instrumental in the final crystallization of these results.
I am especially indebted to Edward Witten for his profound insights and for the invaluable discussions that clarified the correspondence between this twistor-geometric solution and the phenomenological QCD string observed in lattice simulations. His generous engagement with the topological aspects of this work has been a source of significant inspiration.
Declaration of Generative AI in the Writing Process
During the preparation of this work, the author utilized the generative AI Gemini 3 Flash to assist in the technical refinement of the manuscript. Specifically, the tool was employed to:
  • 1.
    Identify and correct typographical errors and grammatical inconsistencies.
  • 2.
    Standardize mathematical notation and ensure the rigorous formatting of LaTeX equations across the text and appendices.
  • 3.
    Improve the clarity and stylistic flow of the scientific narrative while preserving the original physical derivation.
  • 4.
    Cross-verify the consistency of the unified mass-squared formula with experimental data fits and topological sector identifications.
The author has thoroughly reviewed and edited all AI-generated refinements and takes full responsibility for the physical content, mathematical accuracy, and original conclusions presented in this work.
@@
No data was created in this paper. The analytic computations of the Magnus expanstion of the Momentum loop equations were performed in the Mathematica ® notebook published on the Wolfram Cloud[37].